# The Future of Artificial Intelligence and the Mathematical and Physical Sciences (AI+MPS)

### v2

@@ 2.2 Pursue the Science of AI
Across MPS domains, there are significant opportunities for collaboration and progress within the framework of the ``Science of AI.'' Science domain researchers can work together to adapt existing AI tools, leverage scientific frameworks to better understand how and why AI works, and drive the development of new tools that integrate scientific rigor. These efforts would not only help to penetrate the ``black box'' of AI, but would also help to develop AI tools that are embedded from the beginning with scientific principles, making them more robust for scientific applications. By leveraging the deep expertise of the community in uncovering fundamental principles from complex systems, MPS is well-positioned to lead the way in developing AI techniques that are rigorous, interpretable, and firmly rooted in scientific understanding.
@@ 2.2.1 AI Innovations from Science
  • Efficient computing: AI models often require substantial computational resources and large datasets for training. This high demand can lead to inefficiencies, especially when models are applied to complex scientific problems where data may be scarce or expensive to obtain. Often, MPS scientists are using unique state-of-the-art instruments/observatories for which access is extremely limited and competitive. Developing algorithms that are “data efficient” (not just wall time efficient or energy efficient) enables AI to have a meaningful impact on many of the most pressing problems in MPS domains. In general, developing more efficient algorithms is a widely active research area and is critical to improving sustainability as AI is scaled up (see discussion of data efficient methods as a key research opportunity in Sec. 2.5.6). MPS researchers are also contributing to alternative computing platforms (see Sec. 2.8.1).
@@ 2.3 Establish Scalable AI Infrastructures
With the exponentially increasing use of AI, the need for scalable infrastructures is crucial. AI infrastructure comprises the hardware, software, networking resources, and tools to develop, deploy, and manage AI tools and models. The components of AI infrastructure are highly multifaceted, including accessible resources, data management and life-cycle tools, platforms for data discoverability, protocols to make data AI-ready, distributed data systems for caching and connecting data to computing, and software, middleware, and management tools for models (i.e., models-as-infrastructure). Centralizing the infrastructure could increase access, provide support systems, and align expectations across disciplines for better reproducibility (as discussed in Sec. 2.2.3). Existing resources like NSF ACCESS, [^15] https://access-ci.org NSF CloudBank, [^16] https://www.cloudbank.org and the National AI Research Resource (NAIRR) pilot [^17] https://nairrpilot.org are providing a solid foundation to build upon in this regard, but are relatively nascent. Now is the ideal time to identify what the MPS community's needs are so that solutions can be implemented as part of existing plans. This section highlights opportunities to build scalable resources for computing, data management, and curated AI+MPS benchmarks.
@@ 2.3.1 Computing Resources
Increased access and sharing of high-performance computing resources, such as moderate-scale graphics processing unit (GPU) clusters, open-source data, and tool repositories, will help offset the growing costs of AI tools. Practically speaking, compute will increasingly become a utility and it would benefit from being treated as such in the scheme of the broader nationwide infrastructure. Some particularly impactful opportunities for computing resources include:
@@
  • Software development and maintenance: Many essential software tools for AI, such as for automatic differentiation, numerical linear algebra, and numerical equation solvers, have been developed by the MPS community. However, the software and model outcomes of academic research often become inaccessible or inoperable as the personnel who developed them move on to new positions or projects or as funding for development and maintenance ceases. Therefore, there is a need for sustained support for researchers and professionals working on technical software development and code maintenance. One could envision a centralized federal or national infrastructure to provide these services, perhaps as an expansion of NAIRR. This may complement the opportunities and resources available at other federally supported facilities in providing long-term support to make products from federally supported research available to the community and public in the context of a ``models-as-infrastructure'' paradigm. Centralizing these products within a national resource would preserve indefinite access to these tools.
@@
  • Software/hardware interface development: There is opportunity for development at the software/hardware interface for parallelization of ML frameworks on GPUs. While there are many new algorithms for efficient training on small-scale systems, there is less work on practical implementation and inference on large-scale systems, where leadership-class computing is required. For example, massive parallel training on a national cluster is non-trivial. Though it is not very difficult, because we are not at the industry level, there are still barriers to performing mid- to high-tier training. The recent NVIDIA library for equivariant neural networks [^18] https://github.com/NVIDIA/cuEquivariance is a promising step forward, but more focus on this kind of optimization would make these tools more accessible and scalable.
@@ 2.3.2 Data Management and Access
  • Public access to data: Moving toward centralized data that is not held by individual universities would also help democratize AI tools. There are examples of large data activities that have been lost and the data no longer ``exists'' in the public domain. There is a pressing need to make sure that large datasets generated using government funding have long-term storage and support. These need to be curated and freely accessible databases of high quality, structured data (e.g., the Open Reaction Database [^19] https://open-reaction-database.org). Federal support of data structures could use a model similar to the current support of computing centers in ACCESS or NAIRR: operated by partners (usually universities) but with long-term funding by federal sources. Sustainability is a key aspect here; compute resources can be quickly rebuilt when there is a funding lapse but data, once gone, cannot be recovered at scale. Greater access to data could play a role in developing better quality models and lower the cost of using AI. There has been a strong focus on FAIR principles (data should be findable, accessible, interoperable, and reusable) for data, but less so on models. Large amounts of time, resources, and expertise go into constructing models as part of MPS projects, but it is laborious, time consuming, and expensive to maintain and update these models to keep them current and operational. There are few incentives for researchers to do so after expiration of an award or departure of the personnel who created the models. A centralized infrastructure and repository with long-term stable funding and the human expertise to maintain these models and make them available for public access long after expiry of the awards that create them would be valuable in preserving access.
@@ 2.3.3 Benchmarking and Reproducibility
  • Documentation: There is opportunity to develop data hygiene incentives and standards for documenting how AI techniques are used to support knowledge sharing across domains. There are existing tools/best practices that work well enough when the unit tests and benchmarks run on one machine in a few hours, but the large cost of computing and/or data storage required to replicate a training make it impractical to apply standard software engineering tools. To improve reproducibility, researchers could be encouraged to upload the files associated with computation (and associated version of the code).
  • Community standards: Future MPS work will benefit from community-wide agreement on data standards, policies, definitions, and benchmarks. Even though benchmarks are not perfect, they represent a crucial step toward building trust and credibility. This includes establishing best practices for training AI tools and reporting AI findings, which in turn will help standardize data practices and enhance reproducibility. For example, AlphaFold for protein folding and machine-learned density functional theory (DFT) potentials are successful use cases because of structural data standards such as Crystallographic Information Framework (CIF)  [^20] https://www.iucr.org/resources/cif and access to structural databases such as the Inorganic Crystal Structure Database (ICSD)  [^21] https://icsd.products.fiz-karlsruhe.de and the Protein Data Bank (PDB). [^22] https://www.rcsb.org
@@ 2.4 Facilitate Interdisciplinary Collaborations
There are many opportunities for joint AI research projects across MPS domains and with AI experts. Part of what makes collaborations between MPS domains beneficial is that there is often significant overlap in the uses and challenges that integrating AI tools can bring. An advancement in workflow or other solutions in one domain can then be applied to other domains to advance their work, such as the way advances in computer vision enhanced the analysis of scientific images. In addition, fostering collaborations across domains has the potential to eliminate duplicate efforts across subdomains.
@@ 2.4.2 Workshops and Conferences
  • Science at AI conferences: AI conferences could benefit from broadening their scope to include more science applications, as defined by scientists themselves, rather than narrower computer science perspectives. Co-located AI+Science Workshops would also help to raise awareness about science efforts to the AI community. An established example of this is the Machine Learning and the Physical Sciences Workshop, [^23] https://ml4physicalsciences.github.io/2024/ held at the Conference on Neural Information Processing Systems (NeurIPS) for the past several years. Ideally, such workshops would be held regularly, to build a community of researchers who target those workshops to present results.
  • Coding camps/hackathons: Organizing coding camps or hackathons to engage the computer science community on problems that are important to MPS is another way to engage across disciplines. There are already examples of such events that have led to successful collaborations, such as the LHC Olympics anomaly detection data challenge [^24] https://lhco2020.github.io/homepage/ and the Photometric LSST Astronomical Time-Series Classification Challenge (PLAsTiCC) challenge. [^25] https://plasticc.org/ Improved communication between AI and science researchers would enrich both communities and accelerate progress in AI+Science.
@@ 2.5.1 Simulation-Based Inference
Statistical inference is a common task in the MPS domains, where the goal is to determine the underlying parameters of a theoretical model from measured experimental data. To perform classical statistical inference, one needs to know the likelihood, which is the probability of seeing a measured result given the model parameters. Often in MPS, the likelihood is not known exactly, but we can nevertheless use AI to estimate the likelihood from trusted forward simulations that generate synthetic data from a given model. This is the basis for simulation-based inference (SBI), which has been used extensively in the sciences.
@@
The basic framework of SBI is well established, [^26] https://www.pnas.org/doi/10.1073/pnas.1912789117 but there are ample research opportunities to improve the robustness of SBI techniques. For example, the mathematical sciences are developing algorithms for training models to produce simulations for SBI, including ones based on optimal transport. Collaborations between MPS researchers could ensure that these algorithms produced reliable and generalizable results across disciplines.
@@ 2.5.2 Multi-Scale Simulations
AI approaches are revolutionizing multi-scale modeling across MPS disciplines by bridging traditionally separated scales of physical phenomena. In astrophysics, neural networks enable the seamless integration of sub-grid physics models with large-scale cosmological simulations, capturing complex feedback processes that span orders of magnitude in spatial and temporal scales. In materials science and chemistry, AI creates hierarchical models that connect quantum-mechanical calculations at the atomic scale to mesoscale properties and macroscopic behaviors. Physics-informed neural networks (PINNs) and deep operator networks enable the development of surrogate models that preserve physical constraints while drastically reducing computational costs compared to traditional numerical methods.
@@ 2.6.1 AI Co-Pilot
The National Academies identified literature search and complex hypothesis generation as two grand AI challenges for Science. [^27] https://doi.org/10.17226/27457 These are two of the use cases for developing ``AI co-pilots,'' or an interactive AI that assists with performing complex tasks. A key opportunity to support research is to develop AI-based literature search and visualization to instantly return data in tabular, graphical, and/or textual form. This could include AI-based searches that assemble novel databases from across the literature for input into AI-generated code. In addition, synthesis/visualization tools that apply causal AI methods for retrieving structured/unstructured data to create a logically consistent ``worldview'' from the literature corpus can aid scientific discovery. The latter has the potential not just to verify or exclude hypotheses but also to generate ideas for new ones.
@@
The ability of AI to generate plots and synthesize text could be valuable in supporting scientific communication as well. As essentially a ``neutral observer,'' agentic AI can help researchers translate technical content for a broad audience. It can also help to save time with tasks where content is often being repeated, such as reports. Finally, generative reasoning models now allow scientists to create detailed prompts to return code and implementation strategies. As such, AI can speed up hypothesis testing and validation. Using AI in these ways can help researchers to more quickly move on to conducting the research. There is immense interest in applying AI for reasoning tasks in the MPS domains, and this is the first step that could be expanded for additional uses.
@@ 2.8.1 Cost-Efficient Computing
As emphasized in Sec. 2.3, AI requires robust infrastructures, but the energy costs of the compute and data storage resources required to support modern AI-enabled science and technology are non-trivial. As an extreme example, development of GPT-4 is estimated to have cost \$100M to train. [^28] https://www.forbes.com/sites/katharinabuchholz/2024/08/23/the-extreme-cost-of-training-ai-models/ Of course, most applications are not nearly so resource demanding, but as a community we must reckon with the cost-benefit tradeoff in our endeavors. It is important to facilitate access to cutting-edge tools, such as free computing resources for all students and postdocs, but one must also consider the high power cost associated with large-scale computation. How can we capture and price these externalities into AI-enabled science since these efforts have a largely invisible impact on energy? Some potential opportunities to address this include:
@@ 2.8.2 Scientific Integrity
  • Access and ownership: The resources to use AI also need to be broadly accessible to avoid the ability to use AI going to only a privileged few. Continued support for the development and dissemination of resources through the NAIRR Pilot program will be important to continue (see Sec. 2.3.1). AI tools need to be made accessible across the board, with a pathway for software coming out of AI research to be documented and maintained so that it is usable and able to be adapted to the needs of domain scientists. On the other hand, it is important to consider ownership of data used to train the models. There have been many cases where content was used for training purposes without agreement by the owner of the data. Establishing guidelines for the use of data (also see Sec. 2.3.2) that is not clearly open source should be a priority for the MPS community.
@@ 3 Domain-Specific Opportunities
Innovation comes not from the mere use of AI tools, but from the scientific work that AI enables, which could not have been done before, or at least not done as easily. As discussed in Sec. 2.4, a key benefit of facilitating connections and collaborations between AI+MPS researchers is the opportunity for cross-pollination and transferable learning across disciplines. At the same time, it is important to recognize and engage the specific methods, culture, and drivers for each MPS domain, which affects the potential integration and impact of AI in the research process. In this section, we summarize how AI is advancing (and is advanced by) each of the MPS domains and highlight future research opportunities in each area.

### v3

@@ 2.2 Pursue the Science of AI
Across MPS domains, there are significant opportunities for collaboration and progress within the framework of the ``Science of AI.'' Researchers can work together to adapt existing AI tools, leverage scientific frameworks to better understand how and why AI works, and drive the development of new tools that integrate scientific rigor. These efforts would not only help to penetrate the ``black box'' of AI, but would also help to develop AI tools that are embedded from the beginning with scientific principles, making them more robust for scientific applications. By leveraging the deep expertise of the community in uncovering fundamental principles from complex systems, MPS is well-positioned to lead the way in developing AI techniques that are rigorous, interpretable, and firmly rooted in scientific understanding [undefq].
@@ 2.2.1 AI Innovations from Science
  • Efficient computing: AI models often require substantial computational resources and large datasets for training. This high demand can lead to inefficiencies, especially when models are applied to complex scientific problems where data may be scarce or expensive to obtain. Often, MPS scientists are using unique state-of-the-art instruments/observatories for which access is extremely limited and competitive. Developing algorithms that are “data efficient” (not just wall-time efficient or energy efficient) enables AI to have a meaningful impact on many of the most pressing problems in MPS domains. In general, developing more efficient algorithms is a widely active research area and is critical to improving sustainability as AI is scaled up (see discussion of data efficient methods as a key research opportunity in Sec. 2.5.6). MPS researchers are also contributing to alternative computing platforms (see Sec. 2.8.1).
@@ 2.3 Establish Scalable AI Infrastructures
With the exponentially increasing use of AI, the need for scalable infrastructures is crucial. AI infrastructure comprises the hardware, software, networking resources, and tools to develop, deploy, and manage AI tools and models. The components of AI infrastructure are highly multifaceted, including accessible resources, data management and life-cycle tools, platforms for data discoverability, protocols to make data AI-ready, distributed data systems for caching and connecting data to computing, and software, middleware, and management tools for models (i.e., models-as-infrastructure). Centralizing the infrastructure could increase access, provide support systems, and align expectations across disciplines for better reproducibility (as discussed in Sec. 2.2.3). Existing resources like NSF ACCESS [undefi], NSF CloudBank [undefr], and the National AI Research Resource (NAIRR) pilot [undefs] are providing a solid foundation to build upon in this regard, but are relatively nascent. Now is the ideal time to identify what the MPS community's needs are so that solutions can be implemented as part of existing plans. This section highlights opportunities to build scalable resources for computing, data management, and curated AI+MPS benchmarks.
@@ 2.3.1 Computing Resources
Increased access and sharing of high-performance computing resources, such as moderate-scale graphics processing unit (GPU) clusters, open-source and open-weight models, and tool repositories, will help offset the growing costs of AI tools. Practically speaking, compute will increasingly become a utility and it would benefit from being treated as such in the scheme of the broader nationwide infrastructure. Some particularly impactful opportunities for computing resources include:
@@
  • Software development and maintenance: Many essential software tools for AI, such as for automatic differentiation, numerical linear algebra, and numerical equation solvers, have been developed by the MPS community. However, the software and model outcomes of academic research often become inaccessible or inoperable as the personnel who developed them move on to new positions or projects or as funding for development and maintenance ceases. Therefore, there is a need for sustained support for researchers and professionals working on technical software development and code maintenance. This could include hiring Research Software Engineers (RSEs) [undeft], who work with researchers directly to help improve their software. Some labs already invest in RSEs, but one could envision a centralized federal or national infrastructure to provide these services, perhaps as an expansion of NAIRR. This may complement the opportunities and resources available at other federally supported facilities in providing long-term support to make products from federally supported research available to the community and public in the context of a ``models-as-infrastructure'' paradigm. Centralizing these products within a national resource would preserve indefinite access to these tools.
@@
  • Software/hardware interface development: There is opportunity for development at the software/hardware interface for parallelization of ML frameworks on GPUs. While there are many new algorithms for efficient training on small-scale systems, there is less work on practical implementation and inference on large-scale systems, where leadership-class computing is required. For example, massive parallel training on a national cluster is non-trivial. Though it is not very difficult, because we are not at the industry level, there are still barriers to performing mid- to high-tier training. The recent NVIDIA library for equivariant neural networks [undefu] is a promising step forward, but more focus on this kind of optimization would make these tools more accessible and scalable.
@@ 2.3.2 Data Management and Access
  • Public access to data: Moving toward centralized data that is not held by individual universities would also help democratize AI tools. There are examples of large data activities that have been lost and the data no longer ``exists'' in the public domain. There is a pressing need to make sure that large datasets generated using government funding have long-term storage and support. These need to be curated and freely accessible databases of high quality, structured data (e.g., the Open Reaction Database [undefv]). Federal support of data structures could use a model similar to the current support of computing centers in ACCESS or NAIRR: operated by partners (usually universities) but with long-term funding by federal sources. Sustainability is a key aspect here; compute resources can be quickly rebuilt when there is a funding lapse but data, once gone, cannot be recovered at scale. Greater access to data could play a role in developing better quality models and lower the cost of using AI. There has been a strong focus on FAIR principles (data should be findable, accessible, interoperable, and reusable) for data, but less so on models. Large amounts of time, resources, and expertise go into constructing models as part of MPS projects, but it is laborious, time consuming, and expensive to maintain and update these models to keep them current and operational. There are few incentives for researchers to do so after expiration of an award or departure of the personnel who created the models. A centralized infrastructure and repository with long-term stable funding and the human expertise to maintain these models and make them available for public access long after expiry of the awards that create them would be valuable in preserving access.
@@ 2.3.3 Benchmarking and Reproducibility
  • Documentation: There is opportunity to develop data hygiene incentives and standards for documenting how AI techniques are used to support knowledge sharing across domains. There are existing tools/best practices that work well enough when the unit tests and benchmarks run on one machine in a few hours, but the large cost of computing and/or data storage required to replicate a training make it impractical to apply standard software engineering tools. To improve reproducibility, researchers could be encouraged to upload the files associated with computation (and associated version of the code), and/or work with RSEs for benchmark development (see Sec. 2.3.1).
  • Community standards: Future MPS work will benefit from community-wide agreement on data standards, policies, definitions, and benchmarks. Even though benchmarks are not perfect, they represent a crucial step toward building trust and credibility. This includes establishing best practices for training AI tools and reporting AI findings, which in turn will help standardize data practices and enhance reproducibility. For example, AlphaFold for protein folding and machine-learned density functional theory (DFT) potentials are successful use cases because of structural data standards such as Crystallographic Information Framework (CIF) [undefw] and access to structural databases such as the Inorganic Crystal Structure Database (ICSD) [undefx] and the Protein Data Bank (PDB) [undefy].
@@ 2.4 Facilitate Interdisciplinary Collaborations
There are many opportunities for joint AI research projects across MPS domains and with AI experts. Part of what makes collaborations across MPS domains beneficial is that there is often significant overlap in the uses and challenges that integrating AI tools can bring. An advancement in workflow or other solutions in one domain can then be applied to other domains to advance their work, such as the way advances in computer vision enhanced the analysis of scientific images. In addition, fostering collaborations across domains has the potential to eliminate duplicate efforts across subdomains.
@@ 2.4.2 Workshops and Conferences
  • Science at AI conferences: AI conferences could benefit from broadening their scope to include more science applications, as defined by scientists themselves, rather than narrower computer science perspectives. Co-located AI+Science Workshops would also help to raise awareness about science efforts to the AI community. An established example of this is the Machine Learning and the Physical Sciences Workshop [undefz], held at the Conference on Neural Information Processing Systems (NeurIPS) for the past several years. Ideally, such workshops would be held regularly, to build a community of researchers who target those workshops to present results.
  • Coding camps/hackathons: Organizing coding camps or hackathons to engage the computer science community on problems that are important to MPS is another way to engage across disciplines. There are already examples of such events that have led to successful collaborations, such as the LHC Olympics anomaly detection data challenge [undefaa] and the Photometric LSST Astronomical Time-Series Classification Challenge (PLAsTiCC) challenge [undefab]. Improved communication between AI and science researchers would enrich both communities and accelerate progress in AI+Science.
@@ 2.5.1 Simulation-Based Inference
Statistical inference is a common task in the MPS domains, where the goal is to determine the underlying parameters of a theoretical model from measured experimental data. To perform classical statistical inference, one needs to know the likelihood, which is the probability of seeing a measured result given the model parameters. Often in MPS, the likelihood is not known exactly, but we can nevertheless use AI to estimate the likelihood from trusted forward simulations that generate synthetic data from a given model. This is the basis for simulation-based inference (SBI), which has been used extensively in the sciences [undefac].
@@
The basic framework of SBI is well established, but there are ample research opportunities to improve the robustness of SBI techniques [undefad]. For example, the mathematical sciences are developing algorithms for training models to produce simulations for SBI, including ones based on optimal transport. Collaborations between MPS researchers could ensure that these algorithms produced reliable and generalizable results across disciplines.
@@ 2.5.2 Multi-Scale Simulations
AI approaches are revolutionizing multi-scale modeling across MPS disciplines by bridging traditionally separated scales of physical phenomena. In astrophysics, neural networks enable the seamless integration of sub-grid physics models with large-scale cosmological simulations, capturing complex feedback processes that span orders of magnitude in spatial and temporal scales. In materials science and chemistry, AI creates hierarchical models that connect quantum-mechanical calculations at the atomic scale to mesoscale properties and macroscopic behaviors. Physics-informed neural networks (PINNs) and deep operator networks enable the development of surrogate models that preserve physical constraints while drastically reducing computational costs compared to traditional numerical methods [undefae].
@@ 2.6.1 AI Co-Pilot
The National Academies identified literature search and complex hypothesis generation as two grand AI challenges for Science [undefaf]. These are two of the use cases for developing ``AI co-pilots,'' or an interactive AI that assists with performing complex tasks. A key opportunity to support research is to develop AI-based literature search and visualization to instantly return data in tabular, graphical, and/or textual form. This could include AI-based searches that assemble novel databases from across the literature for input into AI-generated code. In addition, synthesis/visualization tools that apply causal AI methods for retrieving structured/unstructured data to create a logically consistent ``worldview'' from the literature corpus can aid scientific discovery. The latter has the potential not just to verify or exclude hypotheses but also to generate ideas for new ones.
@@
The ability of AI to generate plots and synthesize text could be valuable in supporting scientific communication as well. As essentially a ``neutral observer,'' agentic AI can help researchers translate technical content for a broad audience. It can also help to save time with tasks where content is often being repeated, such as reports (like this one). Finally, generative reasoning models now allow scientists to create detailed prompts to return code and implementation strategies. As such, AI can speed up hypothesis testing and validation. Using AI in these ways can help researchers to more quickly move on to conducting the research. There is immense interest in applying AI for reasoning tasks in the MPS domains, and this is the first step that could be expanded for additional uses.
@@ 2.8.1 Cost-Efficient Computing
As emphasized in Sec. 2.3, AI requires robust infrastructures, but the energy costs of the compute and data storage resources required to support modern AI-enabled science and technology are non-trivial. As an extreme example, development of GPT-4 is estimated to have cost \$100M to train [undefag]. Of course, most applications are not nearly so resource demanding, but as a community we must reckon with the cost-benefit tradeoff in our endeavors. It is important to facilitate access to cutting-edge tools, such as free computing resources for all students and postdocs, but one must also consider the high power cost associated with large-scale computation. How can we capture and price these externalities into AI-enabled science since these efforts have a largely invisible impact on energy? Some potential opportunities to address this include:
@@
As discussed in Sec. 2.2.1, developing more efficient algorithms is an active research area in the MPS domains. These MPS efforts can play an important role in addressing the above needs and contribute to energy security and grid modernization in the U.S. by reducing the computational energy costs of AI systems.
@@ 2.8.2 Scientific Integrity
  • Access and ownership: The resources to use AI also need to be broadly accessible to avoid the ability to use AI going to only a privileged few. Continued support for the development and dissemination of resources through the NAIRR Pilot program will be important to continue (see Sec. 2.3.1). AI tools need to be made accessible across the board, with a pathway for software coming out of AI research to be documented and maintained so that it is usable and able to be adapted to the needs of domain scientists. On the other hand, it is important to consider ownership of data used to train the models. There have been many cases where content was used for training purposes without agreement by the owner of the data. Establishing guidelines for the use of data and models that are not clearly open source or open weight should be a priority for the MPS community (see also Sec. 2.3.2).
@@ 3 Domain-Specific Opportunities
Innovation comes not from the mere use of AI tools, but from the scientific work that AI enables, which could not have been done before, or at least not done as easily. As discussed in Sec. 2.4, a key benefit of facilitating connections and collaborations between AI+MPS researchers is the opportunity for cross-pollination and transferable learning across disciplines. At the same time, it is important to recognize and engage the specific methods, culture, and drivers for each MPS domain, which affects the potential integration and impact of AI in the research process. In this section, we summarize how AI is advancing (and is advanced by) each of the MPS domains and highlight future research opportunities in each area. While the discussion below takes a primarily optimistic perspective, we anticipate that substantial original research will be needed to overcome the challenges of integrating AI into each MPS domain.